OpenStudy (anonymous):

Find a solution for the equation. Assume that all angles are acute angles. sec θ = csc(θ + 32°)

OpenStudy (anonymous):

Re-write the left side as 1/cos(θ) and right side as 1/sin(θ + 32°). Now note that cos(θ) = sin(90° - θ). Now we re-write the equation as: 1/sin(90° - θ) = 1/sin(θ + 32°). This implies that 90° - θ = θ + 32°. Re-arranging gives: => 2θ = 58° => θ = 29°

OpenStudy (anonymous):

@scofflawxx